How many liters of a 10% alcohol solution should be mixed with 12 liters of a 20% alcohol solution to abtain a 14% alcohol solution?
Answer by mananth(16949)

qty liters strength%
solution A x-12 10

solution B 12 20

solution C x 14
.1*(x-12)+.2*(12) = .14*x
0.1x-1.2 + 2.4=0.14x
0.04x=1.2
x=1.2/0.04
30 liters
solution A = 18 liters